# Provably Efficient RLHF Pipeline: A Unified View from Contextual Bandits

### v1

@@ Appendix E Proof of Theorem 3
=\ = $\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\widetilde{\theta}_{t}+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}(\theta^{*}-\widetilde{\theta}_{t})$
$+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\widetilde{\theta}_{t}+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}(\theta^{*}-\widetilde{\theta}_{t})$
@@
$+\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a^{\prime}_{t})\big{\rVert}_{\mathcal{H}_{t}^{-1}}+\big{\lVert}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a^{\prime}_{t})\big{\rVert}_{\mathcal{H}_{t}^{-1}}\big{\lVert}\theta^{*}-\widetilde{\theta}_{t}\big{\rVert}_{\mathcal{H}_{t}}$
@@
$+\widetilde{\beta}_{t}\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}+\widetilde{\beta}_{t}\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}$
@@
Here, the first inequality holds since both $a_{t},a^{\prime}_{t}\in\mathcal{A}_{t}$ and by the definition of $\mathcal{A}_{t}$, it implies that
$\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\widetilde{\theta}_{t}$ $\leq\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{\rVert}_{\mathcal{H}_{t}^{-1}},$
$\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a^{\prime}_{t})\big{)}^{\top}\widetilde{\theta}_{t}$ $\leq\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}}.$
The second inequality holds by selection strategy of $(a_{t},a_{t}^{\prime})$, it holds that
$\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\rVert_{\mathcal{H}_{t}^{-1}}$ $\leq\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}},$
$\lVert\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a^{\prime}_{t})\rVert_{\mathcal{H}_{t}^{-1}}$ $\leq\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}.$
@@
${\operatorname*{Reg}}_{T}=\frac{1}{2}\Big{(}\big{(}r(x_{t},\pi^{*}(x_{t}))-r(x_{t},a_{t})\big{)}+\big{(}r(x_{t},\pi^{*}(x_{t}))-r(x_{t},a_{t}^{\prime})\big{)}\Big{)}\leq 2\widetilde{\beta}_{T}\sum_{t=1}^{T}\lVert\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}.$
Furthermore, by the definition of $\mathcal{H}_{t}$, we have
$\mathcal{H}_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}\dot{\sigma}\left(z_{s}^{\top}\widetilde{\theta}_{s+1}\right)z_{s}z_{s}^{\top}\geq\lambda I_{d}+\frac{1}{\kappa}\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}=\frac{1}{\kappa}\left(\kappa\lambda I_{d}+\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}\right)\triangleq\frac{1}{\kappa}V_{t}$
Thus, we obtain
@@
where the first inequality holds by the fact that $\mathcal{H}_{t}\succeq\frac{1}{\kappa}V_{t}$, the second inequality holds by the Cauchy-Schwarz inequality, and the last inequality holds by the elliptic potential lemma in Lemma 4. Thus, we have with probability at least $1-\delta$,
${\operatorname*{Reg}}_{T}\leq 2{\widetilde{\beta}_{T}}\sqrt{2\kappa dT\log\left(1+\frac{4TL^{2}}{\lambda\kappa d}\right)}\leq\widetilde{\mathcal{O}}\left(d\sqrt{{\kappa}{T}}\right).$
This finishes the proof. ∎
@@ Appendix F Supporting Lemmas
Suppose $x_{1},\ldots,x_{t}\in\mathbb{R}^{d}$ and for any $1\leq s\leq t$, $\lVert x_{s}\rVert_{2}\leq L$. Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$. Then, we have
$\sum_{s=1}^{t}\left\|z_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2d\log\left(1+\frac{tL^{2}}{\lambda d}\right).$
@@ G.1 Implementation Details
Datasets.   We use the UltraFeedback-binarized dataset [Rafailov et al., 2023] for the experiments. This dataset is derived from the original UltraFeedback dataset, which comprises 64, 000 prompts sourced from diverse datasets including UltraChat, ShareGPT, Evol-Instruct, TruthfulQA, FalseQA, and FLAN. For each prompt, four model completions were generated using various open-source and proprietary language models, with GPT-4 providing comprehensive evaluations across multiple criteria including helpfulness, honesty, and truthfulness. The binarized version was constructed by selecting the completion with the highest overall score as the ”chosen” response and randomly selecting one of the remaining completions as the ”rejected” response, creating clear preference pairs suitable for reward modeling and direct preference optimization. This dataset structure aligns well with our experimental setup, providing a robust foundation for evaluating different preference learning approaches. The dataset’s diverse prompt sources and evaluation criteria make it particularly valuable for training and evaluating reward models in a real-world context. To further tailor the dataset to our experimental setup, we organize the dataset as follows for the training and deployment stages:
  • Training stage with passive data collection: We randomly sample $T=30,000$ data points from the UltraFeedback-binarized dataset’s train_prefs split for training. Each data point consists of a prompt and two responses with a binary preference label indicating the preferred response. We use the test_prefs split for evaluation.
  • Training stage with active data collection: We allow the method to actively select 6,400 samples from the train_prefs split according to different selection strategies. The global batch size is set to 8 for training. The selection is performed iteratively, where in each iteration, the method selects the most informative samples based on its selection criterion.
  • Deployment stage: We use a pre-processed online variant of the UltraFeedback-binarized dataset from the test_gen split. The dataset is divided into 20 sequential chunks to simulate an online deployment scenario. For each chunk, we generate responses using the current policy (the foundation model of policy model is chosen to be meta-llama / Llama-3.2-1B), evaluate them using both the learned reward model and an oracle reward model. We choose NCSOFT/Llama-3-OffsetBias-RM-8B [Park et al., 2024] as the oracle reward model. After each chunk, we use the policy model to randomly generate 64 responses using different seeds. We then apply various strategies (Random, Best-Two, etc.) to select responses and construct new preference pairs, which are then used to update the reward model and the policy model.
@@ G.2 Validating the Magnitude of $\kappa$
We validate the magnitude of $\kappa$ in Eq. (2) by computing its value during the training process. The results show that $\kappa=171.62\pm 85.49$ during our training process, which is relatively large.
@@ G.3 Combined with Adam Optimizer
Figure 4: For the training stage with passive data collection, we compare our proposed method and MLE [Zhu et al., 2023] in with passive data collection combined with Adam. We report the average accuracy and loss of the reward model during the training process.
G.4 Comparison with DPO
Figure 5: Comparison with DPO in the deployment stage.
We also compare with DPO (Direct Preference Optimization) [Rafailov et al., 2023] in the deployment stage, as a reward-free method, DPO optimizes the policy directly using preference feedback without explicit reward modeling. To ensure a fair comparison, we initialize the policy with 400 samples and use the same dataset settings as PPO to iteratively update the policy model using the DPO algorithm. The results are illustrated in Figure 5. While DPO outperforms the random baseline (Rand-MLE), it achieves lower cumulative rewards than the PromisingSet-based method. This result suggests that DPO’s online learning capability remains a challenge. In contrast, the reward model learned by our selection strategy effectively learned streaming data and continuously updates the policy as new data arrive, indicating that in our deployment stage, a reward model with PPO may be a more suitable choice for sequentially learning from new data.

### v2

@@ Appendix E Proof of Theorem 3
=\ = $2\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\theta^{*}+\big{(}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\theta^{*}$
=\ = $2\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}(\theta^{*}-\widetilde{\theta}_{t})+2\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t})\big{)}^{\top}\widetilde{\theta}_{t}$
$+\big{(}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}(\theta^{*}-\widetilde{\theta}_{t})+\big{(}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\widetilde{\theta}_{t}$
@@
$+\big{\lVert}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}}\big{\lVert}\theta^{*}-\widetilde{\theta}_{t}\big{\rVert}_{\mathcal{H}_{t}}+\big{(}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\widetilde{\theta}_{t}$
@@
$+\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}}$
@@
$+\big{(}\phi(x_{t},\pi^{*}(x_{t}))-\phi(x_{t},a_{t}^{\prime})\big{)}^{\top}\widetilde{\theta}_{t}+\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}}$
=\ = $3\widetilde{\beta}_{t}\big{\lVert}\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}},$
where the first inequality holds by the Holder’s inequality and the arm selection strategy of $a_{t}$ such that $\phi(x_{t},\pi^{*}(x_{t}))^{\top}\widetilde{\theta}_{t}\leq\phi(x_{t},a_{t})^{\top}\widetilde{\theta}_{t}$, the second inequality holds by $\widetilde{\theta}_{t}\in\mathcal{C}_{t}$ with probability at least $1-\delta$ by Lemma 1, the third inequality holds by arm selection strategy of $a_{t}^{\prime}$ such that $a_{t}^{\prime}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(x_{t},a)^{\top}\widetilde{\theta}_{t}+2\widetilde{\beta}\lVert\phi(x_{t},a)-\phi(x_{t},a_{t})\rVert_{\mathcal{H}_{t}^{-1}}$.
By the definition of $\mathcal{H}_{t}$, we have
$\mathcal{H}_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}\dot{\sigma}\left(z_{s}^{\top}\widetilde{\theta}_{s+1}\right)z_{s}z_{s}^{\top}\geq\lambda I_{d}+\frac{1}{\kappa}\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}=\frac{1}{\kappa}\left(\kappa\lambda I_{d}+\sum_{s=1}^{t-1}z_{s}z_{s}^{\top}\right)=\frac{1}{\kappa}V_{t}.$
@@
where the first inequality holds by the fact that $\mathcal{H}_{t}\succeq\frac{1}{\kappa}V_{t}$, the second inequality holds by the Cauchy-Schwarz inequality, and the last inequality holds by the elliptic potential lemma in Lemma 4.
Therefore, we have
${\operatorname*{Reg}}_{T}\leq\frac{3}{2}{\widetilde{\beta}_{T}}\sqrt{2\kappa dT\log\left(1+\frac{4\kappa TL^{2}}{\lambda d}\right)}\leq\widetilde{\mathcal{O}}\big{(}d\sqrt{\kappa T}\big{)}.$
where the This completes the proof. ∎
@@ Appendix F Supporting Lemmas
Suppose $x_{1},\ldots,x_{t}\in\mathbb{R}^{d}$ and for any $1\leq s\leq t$, $\lVert x_{s}\rVert_{2}\leq L$. Let $V_{t}=\lambda I_{d}+\sum_{s=1}^{t-1}x_{s}x_{s}^{\top}$ for $\lambda\geq 0$. Then, we have $\sum_{s=1}^{t}\left\|z_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2d\log\left(1+\frac{tL^{2}}{\lambda d}\right)$.
@@ G.1 Implementation Details
Datasets.   We use the UltraFeedback-binarized dataset [Rafailov et al., 2023] for the experiments. This dataset is derived from the original UltraFeedback dataset, which comprises 64, 000 prompts sourced from diverse datasets including UltraChat, ShareGPT, Evol-Instruct, TruthfulQA, FalseQA, and FLAN. For each prompt, four model completions were generated using various open-source and proprietary language models, with GPT-4 providing comprehensive evaluations across multiple criteria including helpfulness, honesty, and truthfulness. The binarized version was constructed by selecting the completion with the highest overall score as the "chosen" response and randomly selecting one of the remaining completions as the "rejected" response, creating clear preference pairs suitable for reward modeling and direct preference optimization. This dataset structure aligns well with our experimental setup, providing a robust foundation for evaluating different preference learning approaches. The dataset’s diverse prompt sources and evaluation criteria make it particularly valuable for training and evaluating reward models in a real-world context. To further tailor the dataset to our experimental setup, we organize the dataset as follows:
  • Online RLHF with passive data collection: We randomly sample $T=30,000$ data points from the UltraFeedback-binarized dataset’s train_prefs split for training. Each data point consists of a prompt and two responses with a binary preference label indicating the preferred response. We use the test_prefs split for evaluation.
  • Online RLHF with active data collection: We allow the method to actively select 6,400 samples from the train_prefs split according to different selection strategies. The global batch size is set to 8 for training. The selection is performed iteratively, where in each iteration, the method selects the most informative samples based on its selection criterion.
  • Deployment-time adaption: We use a pre-processed online variant of the UltraFeedback-binarized dataset from the test_gen split. The dataset is divided into 20 sequential chunks to simulate an online deployment scenario. For each chunk, we generate responses using the current policy (the foundation model of policy model is chosen to be meta-llama / Llama-3.2-1B), evaluate them using both the learned reward model and an oracle reward model. We choose NCSOFT/Llama-3-OffsetBias-RM-8B [Park et al., 2024] as the oracle reward model. After each chunk, we use the policy model to randomly generate 64 responses using different seeds. We then apply various strategies (Random, Best-Two, etc.) to select responses and construct new preference pairs, which are then used to update the reward model and the policy model.
@@ G.2 Validating the Magnitude of $\kappa$
We validate the magnitude of $\kappa$ by computing its value during the training process. The results show that $\kappa=171.62\pm 85.49$ during our training process, which is relatively large.
@@ G.3 Combined with Adam Optimizer
G.4 Comparison with DPO
We also compare with DPO [Rafailov et al., 2023] in the deployment stage, as a reward-free method, DPO optimizes the policy directly using preference feedback without explicit reward modeling. To ensure a fair comparison, we initialize the policy with 400 samples and use the same dataset settings as PPO to iteratively update the policy model using the DPO algorithm. The results are illustrated in Figure 11. While DPO outperforms the random baseline (Rand-MLE), it achieves lower cumulative rewards than the methods using our action selection. This result suggests that DPO’s online learning capability remains a challenge. In contrast, the reward model learned by our selection strategy effectively learned streaming data and continuously updates the policy as new data arrive, indicating that in our deployment stage, a reward model with PPO may be a more suitable choice for sequentially learning from new data.
@@ G.4 Comparison with DPO
Figure 9: For online RLHF with passive data collection, we compare our proposed method and MLE [Zhu et al., 2023] in with passive data collection combined with Adam. We report the average accuracy and loss of the reward model during the training process.
Figure 10: Comparison of DPO and our method in deployment-time adaptation.
Figure 11: Comparison of different methods for full update in deployment-time adaptation.
G.5 Full Update of Reward Model
Figure 11 shows deployment-time adaptation results using the meta-llama/Llama-3.2-1B model, where we update all parameters of the reward model instead of only the final layer. Both our method and MLE use the same action selection strategy. Our approach achieves comparable performance with MLE, indicating that our OMD-based method is still compatible with full-model updates.